4-30 A man wants to help provide a college education for A his young daughter. He can afford to invest $600/yr for the next 4 years, beginning on the girls 4th birthday. He wishes to give his daughter $4000 on her 18th, 1gth, 20th, and 21st birthdays, for a total of $16,000. Assuming 5% interest, what uniform annual investment will he have to make on the girls 8th through 17th birthdays? Tameshia deposits $5500 in her retirement account every year. If her account pays an average of 6% interest and she makes 38 deposits before she retires, how much money can she withdraw in 20 equal payments beginning one year after her last deposit? Contributed by Paul R. McCrigh, University of 4-27 South Florida 4-23 Jennifer Creek is saving up for a new car. She wa nts 10,000 of the $26,000 esti- years. She deposits $5000 into a ill make monthly deposits the next two years. If the savings account pays a monthly compounding, how much must she deposit each month? Contributed by Gillian Nicholls, Southeast to finance no more than $ mated price in two savings account now and w for nominal interest rate of 3% per year with uri State Universiy
Answer #1

4-27

Annual deposit = $5500

Number of deposits = 38

R = 6%

So,

Future value of deposits = 5500*(1.06^38 - 1)/.06

Future value of deposits = $747473.1

Let, annual withdrawals = P

No. of withdrawals = 20

Then,

747473.1 = P*(1-1/1.06^20)/.06

P = 747473.1/((1-1/1.06^20)/.06)

P = $65168.11 or $65168

So, value of annual 20 withdrawals will be $65168.

Pl. repost other unanswered questions for their proper answers!